the complement of an angle is 20 degrees less than half the supplement find the angle its complement and its supplement
90-x=(180-x)/2 -20

Answer by gonzo(654)

you have the formula right.
all you have to do is solve it.
-----
90 - x = (180-x)/2 - 20
multiply both sides of equation by 2.
180 - 2*x = 180 - x - 40
subtract 180 from both sides and add 2*x to both sides and add 40 to both sides.
equation becomes
40 = x which is the same as x = 40.
-----
complement of 40 is 50.
supplement of 40 is 140.
1/2 * 140 - 20 should equal 50.
70 - 20 = 50
it does.
formula is good.
angle equals 40 degrees.